### v1

# Long-term prediction of El Niño-Southern Oscillation using reservoir computing with data-driven realtime filter

Abstract
In recent years, the application of machine learning approaches to time-series forecasting of climate dynamical phenomena has become increasingly active. It is known that applying a band-pass filter to a time-series data is a key to obtaining a high-quality data-driven model. Here, to obtain longer-term predictability of machine learning models, we introduce a new type of band-pass filter. It can be applied to realtime operational prediction workflows since it relies solely on past time series. We combine the filter with reservoir computing, which is a machine-learning technique that employs a data-driven dynamical system. As an application, we predict the multi-year dynamics of the El Niño-Southern Oscillation with the prediction horizon of 24 months using only past time series.
preprint:
Lead paragraph: Predicting long-term climate phenomena, such as the El Niño–Southern Oscillation (ENSO), is crucial for mitigating the impacts of extreme weather events worldwide on agriculture and economics. Yet, forecasting the complex systems remains a challenge, especially when different time scales of variability overlap. Furthermore, in past studies which use convectional smoothing technique, information from the future is introduced to some extent. In this paper, we introduce a new type of filter that is applicable to realtime operational forecast and combine it with a machine-learning framework known as reservoir computing, a brain-inspired method for dynamical systems modeling. An important strength of our method is that it relies solely on past data, enabling realtime predictions without the need for future observations. The parameters of the realtime filter as well as the hyperparameters of the resevoir computing model are systematically calibrated with the Bayesian optimizations. Using this method, we show that the prediction horizon of ENSO can be extended up to two years, which is longer than most of prediction horizons previously reported. While we focus on ENSO forecasts in this article, our method can also be applied to a wide range of complex, high-dimensional phenomena that demand robust and efficient long-term forecasting.
I Introduction
The El Niño-Southern Oscillation (ENSO) is the primary interannual fluctuation in the Earth’s climate. It is characterized by the temperature anomaly in the central to eastern tropical Pacific Ocean. ENSO shows an oscillation among warm (El Niño), cold (La Niña), and neutral phases with a typical timescale of 3 to 8 years Bjerknes (1969). With the nature of teleconnection, climate variability has a strong effect on global atmospheric circulation, ecological systems, public health, and economics McPhaden, Zebiak, and Glantz (2006).
The prediction of the time evolution of ENSO has been continuously addressed throughout the past decades. The approaches include conceptual models Schopf and Suarez (1990); Battisti and Hirst (1989); Wyrtki (1985); Cane, Münnich, and Zebiak (1990), statistical models Trenberth and Hoar (1996); Grieger and Latif (1994), and global climate models (GCMs)Luo et al. (2008); Guilyardi et al. (2009). The prediction using fully-coupled ocean-atmosphere models has demonstrated the high potential for ENSO prediction accuracy, yet challenges remain for long-term forecasts beyond a one-year lead time Tang et al. (2018).
In recent years, as with the prediction of tropical cyclones Loi, Wu, and Liang (2024), there has been a notable improvement in the prediction skills of ENSO using machine learning as well as using methods of complexity science Ludescher et al. (2013); Meng et al. (2020). There are historical examples of using neural networks to model nonlinear dynamics of ENSO Grieger and Latif (1994); Timmermann, Voss, and Pasmanter (2001); Guardamagna et al. (2024). In this century, it has become possible to achieve forecasts exceeding one year, with performance comparable to that of GCMs Ham, Kim, and Luo (2019); Chen, Gilani, and Harlim (2021); Chen et al. (2023); Wang et al. (2024).
The reservoir computing is a brain-inspired machine-learning technique that employs a data-driven dynamical system Jaeger (2001); Jaeger and Haas (2004); Lu et al. (2017); Lu, Hunt, and Ott (2018). It is a method effective in predicting time series and frequency spectra in chaotic behaviors and is capable of learning time series of high-dimensional dynamical systems such as fluid flow  Nakai and Saiki (2018, 2021); Kobayashi et al. (2021); Nakai and Saiki (2024). In particular, Nakai and Saiki Nakai and Saiki (2021) clarified that delay-coordinates with appropriate delay time and dimension of the delay-coordinates are efficient when the number of observable variables is smaller than the effective dimension of the attractor. The reservoir computing can predict the statistical quantity of a particular macroscopic variable of chaotic fluid flow, which cannot be calculated from a direct numerical simulation of the Navier–Stokes equation because of its high computational cost Kobayashi et al. (2021). The reservoir computing approach has been proven to be effective in prediction of complex atmospheric systems such as Madden-Julian Oscillation Suematsu et al. (2022), the Asian summer monsoon Mitsui and Boers (2021), North Atlantic Oscillation Huang and Fu (2023) as well as ENSO Hassanibesheli, Kurths, and Boers (2022); Guardamagna et al. (2024).
The aim of this study is to establish a general methodology for performing long-term prediction of various time series in operational realtime formulation, which excludes information from the future. Applying the band-pass filter to the time series data is proven to be effective for extending the prediction horizon of the ENSO time series. For a robust long-term prediction, we introduce a new data-driven filter. It is known that the high-frequency disturbances degrade the performance of the machine learning prediction, while the low-frequency behavior such as trends and multidecadal variability make construction of machine-learning model difficult Hassanibesheli, Kurths, and Boers (2022). In past studies, the formulation of predicting ENSO by artificial neural networks requires the application of some convectional bandpass filter techniques such as moving averageGuardamagna et al. (2024) or Butterworth filter Hassanibesheli, Kurths, and Boers (2022); Ham, Kim, and Luo (2019), which incorporate information from the future time steps into the filtered time series. Strictly speaking, even the convectional Niño indices include future information by the moving average. Therefore, those methodologies are not practical to be applied in a realtime operational forecast. This study is the first successful attempt to develop a filtering method that extracts signals with cycles of 3 to 8 years using only past time series. For the filtered Niño-3.4 time series data, we construct a data-driven model. In this study, we employ reservoir computing to construct a machine-learning model capable of predicting the multi-year dynamics of the ENSO for over two years.
II Data and Methods
II.1 Realtime fitering method
In this subsection, we present a new method to find a suitable filter for the corresponding time series data. Note that the corresponding Python codes of realtime filter and reservoir computing combining with Bayesian optimization are available in the repository Yoneda (2024). The pure reservoir computing code is also available.Tanaka, Nakane, and Hirose (2021) To find the appropriate filter, we use Bayesian optimization and the following weight function of the moving average:
$\Psi(t)=\underbrace{\left(d_{1}\cos\left(\frac{t}{\pi r_{1}}\right)+d_{2}\cos\left(\frac{t}{\pi r_{2}}\right)\right)}_{passing\ freq}\underbrace{\frac{(w-t)^{c}}{w^{c}}}_{band\ width}\quad\text{for}\quad t\in[0,w],$
where $d_{1},d_{2},r_{1},r_{2},c,w>0$ are the parameters of the corresponding objective function (3). Let $\Omega\subset\mathbb{Z}$ be a prescribed finite sequence. Having obtained $\Psi$ by the optimization, then we applied the weighted moving average to the original data $y:\Omega\to\mathbb{R}$ as follows:
[…]
We emphasize that since the weight function $\Psi$ is localized in the timeline, in particular, it does not contain any future information, this bandwidth cannot have a unique threshold due to the Fourier uncertainty principle. Here is a rough explanation of the bandwidth. If $c$ is close to zero, then the Fourier transform of the corresponding weight function is close to the sinc function, which represents narrow bandwidth (compared to the case when $c$ is large). To the contrary, if $c$ tends to infinity, then the Fourier transform of the corresponding weight function tends to identically one (since $\Psi$ tends to the Dirac-delta function), which represents broad bandwidth (compared to the case when $c$ is small). The main philosophy of this filter method is to effectively reduce the “indefinite factor” (such as white noise) without using the conventional Fourier transform, since the time domain of the training data $\Omega$ is, of course, always finite, and, applying the Fourier transform to such finite elements is no longer mathematically rigorous. To effectively reduce the indefinite factor, we use the idea of “dictionary”, in other word, “key-value pairs”. First, we discretize the range of the filtered data $y^{*}$ as follows: For an integer $K$ which is greater than $1$, we choose $\{a_{k}\}_{k=1}^{K}\subset\mathbb{R}$ such that $a_{1}<a_{2}<\cdots<a_{K}$.
Figure 1: Optimization of filter performances. a) A filtered time series data $y^{*}$ (black thick line) is classified into discrete values $\tilde{y}$ (blue dots). b) Concept of the key and value: The indefinite factor of the filter is reduced statistically by maximizing (3), which represents the highest rate at which a key pattern $\sigma^{L}_{n}$ is followed by one of the values $\{a_{k}\}^{K}_{k=1}$ at the subsequent time step.
Remark II.1
To determine the positions of $\{a_{k}\}_{k=1}^{K}$, we have applied the cumulative distribution function to $y^{*}$, more precisely, #$\#\{t\in\Omega:a_{k}\leq y^{*}(t)<a_{k+1}\}$ ($k=1,2,\cdots,K-1$), #$\#\{t\in\Omega:y^{*}(t)<a_{1}\}$ and #$\#\{t\in\Omega:a_{K}\leq y^{*}(t)\}$ to be almost the same integer.
Using this $\{a_{k}\}_{k=1}^{K}$, we classify the filtered data $y^{*}$ of each function value as follows:
$\tilde{y}(t):=\text{argmin}_{a\in\{a_{k}\}_{k=1}^{K}}|y^{*}(t)-a|\quad\text{for}\quad t\in\Omega.$ (1)
See Fig.1 a. To perform the pattern classification, we first choose the maximum and minimum length of the pattern a-priori: $L_{max}$ and $L_{min}$ ($L_{max}>L_{min}\geq 1$), and let $\sigma_{n}^{L}$ ($n=1,2,\cdots,N_{L}$) be a permutation operator (we say “key”) such that
$\sigma_{n}^{L}:\{1,2,\cdots,L\}\to\{a_{1},a_{2},\cdots,a_{K-1},a_{K}\}\qquad(\ell\mapsto\sigma_{n}^{L}(\ell))$
for $L\in[L_{min},L_{max}]\cap\mathbb{Z}$ with $\sigma_{n}^{L}\not=\sigma_{n^{\prime}}^{L}$ ($n\not=n^{\prime}$). Then there exists a suitable $N_{L}$ such that the permutation operators $\{\sigma^{L}_{n}\}_{n=1}^{N_{L}}$ satisfy the following two properties:
$\begin{cases}\text{For any $t\in\Omega$, there is $n\in\{1,\cdots,N_{L}\}$ such that}\\ \qquad\text{$\sigma_{n}^{L}(\ell)=\tilde{y}(t-\ell)$ for $\ell=1,2,\cdots,L$},\\ \text{For any $n\in\{1,\cdots,N_{L}\}$ there is $t\in\Omega$ such that}\\ \qquad\text{$\sigma_{n}^{L}(\ell)=\tilde{y}(t-\ell)$ for $\ell=1,2,\cdots,L$.}\end{cases}$ (2)
Note that $N_{L}\leq K^{L}$ due to the sequence with repetition. We now define sets of discrete-time $\Sigma_{n}^{L}\subset\Omega$ such that
$\Sigma_{n}^{L}:=\{t\in\Omega:\sigma_{n}^{L}(\ell)=\tilde{y}(t-\ell)\quad\text{for}\quad\ell=1,2,\cdots,L\}.$
Then we can reduce the indefinite factor by using $\{\tilde{y}(t)\}_{t\in\Sigma_{n}^{L}}$ which is called “value”, as displayed schematically in Fig.1 b). Let
$\Pi_{n}^{L}(k):=\{t\in\Omega:\sigma_{n}^{L}(\ell)=\tilde{y}(t-\ell)\quad\text{for}\quad\ell=1,2,\cdots,L\quad\text{and}\quad\tilde{y}(t)=a_{k}\},$
and we maximize the following objective function under the condition that the autocorrelation between the original and the filterd data is not small:
$\frac{\displaystyle\sum_{L=L_{min}}^{L_{max}}\#\left\{n:\text{there is a }k\in\{1,2,\cdots,K\}\text{ such that }\frac{\#\Pi^{L}_{n}(k)}{\#\Sigma_{n}^{L}}\geq\gamma\right\}}{\displaystyle\sum_{L=L_{min}}^{L_{max}}N_{L}},$ (3)
where $\gamma\in(0,1]$ is the match rate of $a_{k}$ for each $\sigma_{n}^{L}$. Here, $\gamma$, as well as $L_{max}$ and $L_{min}$ are prescribed parameters in the numerical implementation. $\gamma$ is close to 1. $L_{max}$ and $L_{min}$ are determined so that the length of frequent patterns are sufficiently covered.
II.2 Data description
We use monthly sea surface temperature (SST) for the 1870–2022 period, retrieved from the Hadley Centre Global Sea Ice and Sea Surface Temperature dataset Centre (2006). The one-degree latitude-longitude grid data is averaged over the $5^{\circ}$–$5^{\circ}$, $170^{\circ}$–$5^{\circ}$ regions in the tropical central and eastern Pacific, which are commonly used to define the Niño-3.4 index. To obtain the SST anomaly from the typical seasonal cycle, the 1971–2000 monthly climatological values are subtracted. We then filter the unrequired frequency ranges by the realtime filter described in the previous subsection. The convectional Niño-3.4 index uses 5-month running mean to define El Niño and La Niña events. However, in this study, we utilize monthly SST values without applying any moving average to exclude future information and enable realtime prediction. Hereafter, we refer to this filtered time series as the realtime SST anomaly.
Figure 2: Schematic diagram of reservoir computing. In training phase ($t<t_{0}$), optimal value of $\mathbf{W}_{\text{out}}$ is determined using the input time series with prescribed random matrix $\mathbf{A}$ and $\mathbf{W}_{\text{in}}$. In the prediction phase ($t\geq t_{0}$), the dynamics of the filtered time series are inferred recursively by the optimized model of the output matrix, which is denoted as $\mathbf{W}^{*}_{\text{out}}$.
II.3 Reservoir computing
The echo state network is a widely used reservoir computing architecture based on a recurrent neural network, consisting of an input layer, a reservoir state vector, and an output layerJaeger (2001); Jaeger and Haas (2004); Lu et al. (2017); Lu, Hunt, and Ott (2018). A vector of an observed variable as input and a reservoir state vector at time step $t$ is updated as
$\begin{cases}\mathbf{r}(t+\Delta t)=(1-\alpha)\mathbf{r}(t)+\alpha\tanh(\mathbf{A}\mathbf{r}(t)+\sigma_{\text{in}}\mathbf{W}_{\text{in}}\mathbf{u}(t))\\ \mathbf{u}(t+\Delta t)=\mathbf{W}_{\text{out}}\mathbf{r}(t+\Delta t),\end{cases}$ (4)
where $\mathbf{u}(t)\in\mathbb{R}^{M}$ is the input vector of the time series data, $\mathbf{r}(t)\in\mathbb{R}^{N}~{}(N\gg M)$ is the reservoir state vector, $\Delta t$ represents the time interval at which the time series is updated (1 month in this study), and $\sigma_{\text{in}}$ is the scaling parameter of the input matrix $W_{\text{in}}$. $\mathbf{W}_{\text{in}}\in\mathbb{R}^{N\times M}$ and $\mathbf{W}_{\text{out}}\in\mathbb{R}^{M\times N}$ are input and output matrices; $\mathbf{A}\in\mathbb{R}^{N\times N}$ is a matrix that depicts the recurrent connectivity within the reservoir. The density of the non-zero elements of $\mathbf{A}$ is determined by the parameter $p$. The maximum eigenvalue of $\mathbf{A}$ is controlled by the parameter $\rho$ $(|\rho|<1)$ to satisfy the echo state property that affects the performance of the data-driven model. $\alpha$ ($0<\alpha\leq 1$) is the coefficient called leaking rate that adjusts the smoothness of the time evolution of $\mathbf{r}(t)$. Hyperbolic tangent function for a vector $\mathbf{q}=(q_{1},q_{2},\ldots,q_{N})^{\text{T}}$ is defined as $\tanh(\mathbf{q})=(\tanh(q_{1}),\tanh(q_{2}),\ldots,\tanh(q_{N}))^{\text{T}}$ where superscript T represents the transpose.
The choice of input and output variables of the reservoir computing and their dimensions is essential for constructing the modeling. It is known that a chaotic attractor can be reconstructed by some observable variable and its delay-coordinates Takens (1981); Sauer and Yorke (1991). For the reservoir computing models, it has been reported that a delay-coordinate variable with appropriate delay time is efficient when the number of observable variables is smaller than the effective dimension of the attractor Nakai and Saiki (2021); Huang, Yang, and Fu (2020); Huang and Fu (2023). Thus, we employ the choice of delay coordinates of an observable variable as a hyperparameter of the model.
Let $y^{*}$ be the realtime SST anomaly. For these reasons above, we introduce an $M$-dimensional delay-coordinate vector of realtime SST anomaly with a delay-time $\Delta\tau$ as input and output data $\mathbf{u}$ in Eq. (4):
$\mathbf{u}(t)=(y^{*}(t),y^{*}(t-\Delta\tau),\ldots,y^{*}(t-(M-1)\Delta\tau))^{T},$
where the optimal values of $(\Delta\tau,M)$ are determined using the Bayesian optimization framework Optuna Akiba et al. (2019), along with other hyperparameters as described in the next section. For each set of training and prediction, let $t_{0}$ be the initial time step of prediction. The preceding 100-year period of the data $\mathbf{u}(t)\ (t_{0}-1200\leq t\leq t_{0}-1)$ is used as training data.
In the training phase, with a fixed pair of matrices $\mathbf{A}$ and $\mathbf{W}_{\text{in}}$, $\mathbf{W}^{*}_{\text{out}}$ is optimized using training data so that $\mathbf{W}^{*}_{\text{out}}\mathbf{r}(t)\approx\mathbf{u}(t)$ is satisfied (Fig.2 left). Here the cost function
$\Lambda=\sum^{T_{\text{train}}}_{t=1}||\mathbf{W}_{\text{out}}\mathbf{r}(t)-\mathbf{u}(t)||+\beta||\mathbf{W}_{\text{out}}||^{2}$ (5)
is minimized to obtain the best inference of the model, where $T_{\text{train}}$ is the length of training data and ||..||||..||| | . . | | is $L_{2}$-norm of a vector. The optimal output matrix is
$\mathbf{W}^{*}_{\text{out}}=\delta\mathbf{U}\delta\mathbf{R}^{T}(\delta\mathbf{R}\delta\mathbf{R}^{T}+\beta\mathbf{I})^{-1}$ (6)
where the columns of the matrices $\delta\mathbf{R}$ and $\delta\mathbf{U}$ corresponds to the vectors $\mathbf{r}(t)$ and $\mathbf{u}(t)$, respectively, in chronological order. $\mathbf{I}$ denotes identity matrix.
In the prediction phase, the output matrix is fixed to the optimal value $\mathbf{W}^{*}_{\text{out}}$ obtained (Fig. 2 right). We then get the predicted value $\hat{\mathbf{u}}(t+\Delta t)$ from $\mathbf{u}(t)$ and $\mathbf{r}(t)$, using Eq. (4) with the same $\mathbf{A}$ and $\mathbf{W}_{\text{in}}$. The output vector is fed back as input to predict the value at the next time step $t+2\Delta t$. In this manner, the time integration is recursively advanced to obtain the time evolution up to 36 months ahead.
II.4 Bayesian optimization of hyperparameters
For both the realtime filter and the reservoir model, the combinations of hyperparameters are selected using Bayesian optimization with Optuna Akiba et al. (2019). The hyperparameters of the realtime filter are chosen to maximize the product of the objective function in Eq. (3) when applied to the monthly SST anomaly, and the maximum lag-correlation between the original and filtered monthly SST anomaly. The hyperparameters of the reservoir computing model, namely, $M$, $\Delta\tau$, $N$, $\beta$, $p$, $\rho$, $\alpha$, and $\sigma_{\text{in}}$, as well as the random matrices $\mathbf{A}$ and $\mathbf{W_{\text{in}}}$, are chosen based on the prediction performance of 120 sequences of training and predictions whose $t_{0}$ ranges from January 1986 to December 1995. As a measure of the model performance, we use the all-season correlation skill $C(\mu)$
$C(\mu)=\sum_{m=1}^{12}\frac{\sum_{\eta=\eta_{s}}^{\eta_{e}}(Y_{\eta,m}-\overline{Y_{m}})(P_{\eta,m,\mu}-\overline{P_{m,\mu}})}{\sqrt{\sum_{\eta=\eta_{s}}^{\eta_{e}}(Y_{\eta,m}-\overline{Y_{m}})^{2}\sum_{\eta=\eta_{s}}^{\eta_{e}}(P_{\eta,m,\mu}-\overline{P_{m,\mu}})^{2}}}$ (7)
as a function of the forecast lead months $\mu$. Here, $Y$ and $P$ denote the observed and the predicted values, respectively. $\overline{Y_{m}}$ and $\overline{P_{m,l}}$ denote the climatologies with respect to the calendar month $m$ (from 1 to 12) and the forecast lead months $\mu$. The label $\eta$ denotes the forecast target year. $\eta_{s}$ and $\eta_{e}$ denote the earliest (that is, 1986) and the latest year (that is, 1995) of the period, respectively. $C(\mu=24)$ is maximized here. The determination of the random matrices $\mathbf{W}_{\text{in}}$ and $\mathbf{A}$ is as follows. First, using two consecutive integers $i$ and $i+1$ as seed values, two $N\times N$ matrices $\mathbf{W}_{i}$ and $\mathbf{W}_{i+1}$ are generated with a density of non-zero elements equal to $p$. Next, by linearly interpolating between the two matrices while maintaining the density $p$, a matrix $\mathbf{W}_{i+x}$ characterized by a real seed value $x$ $(0<x<1)$ is created. Finally, $\mathbf{W}_{i+x}$ is scaled to match the condition of spectral radius $\rho$.
III Result
III.1 Performance of the new filter
As a result of the Bayesian optimization, the best set of filter parameters is obtained as shown in Table 1. The shape of the filter function is displayed in Fig.3a. The weight equals zero everywhere $t\geq 0$. The frequency response of the filter. The filter is applied by performing a convolution integral of the weighting function shown in Fig.3a with the original time series. Fig.3b shows that the passband of the filter lies in about four to eight-year periods, which is roughly equal to the typical timescale of the perturbation of ENSO.
Parameter Value
$r_{1}$ 39.333
$r_{2}$ 2.789
$d_{1}$ 0.152
$d_{2}$ 0.448
$c$ 1.086
$w$ 65
Table 1: The list of parameters and their values for the realtime filter obtained from the Bayesian optimization.
Figure 3: Characteristics of the new realtime filter. (a) The weight function in terms of time step and (b) frequency response function of the weight function. The weight function for $t\geq 0$ is zero so that the filtered time series does not include information from the future. The left endpoint of the weight function is also determined by Optuna in the parameter $w$.
By applying the new realtime filter, we obtain a target time series that preserves information on long-term ENSO variations while enhancing the predictive performance of the reservoir computing model. The filtered realtime SST anomaly is shown in Fig.4a along with the original time series. Due to the asymmetry of the filter function with respect to the positive and negative time ranges, the overall time series shifts backward by approximately five months. The shift is highlighted by the lag correlation between the original and the filtered time series (Fig.4b). The maximum correlation is 0.837 with a lag of five months, and we assume that the filtered realtime SST anomaly is can be used as an alternative index for the state of ENSO.
Figure 4: Comparison between the original and filtered realtime SST anomaly. (a) The time series of original (black) and filtered (red) indices from 1980 to 2022. (b) Lag correlation between the two time series over the entire data period. When the lag $\lambda$ is positive, it corresponds to delaying the original time series by $\lambda$ steps.
III.2 Prediction skill of the model
As a result of the Bayesian optimization, the optimal set of the hyper parameters for reservoir computing model is obtained as listed in Table 2. To evaluate the prediction skill of the current method, the sequence of training and prediction is conducted 180 times with different $t_{0}$ every month between January 2001 and December 2015. The forecast skill of the filtered realtime SST anomaly was evaluated using the all-season correlation skill in Eq. (7) with $\eta_{s}=2001$ and $\eta_{e}=2015$.
Parameter Value
$\Delta\tau$ 4
$M$ 9
$N$ 244
$\beta$ 0.759
$p$ 0.290
$\sigma_{\text{in}}$ 0.477
$\rho$ 0.712
$\alpha$ 0.975
Table 2: The list of parameters and their values for the reservoir computing model obtained from the Bayesian optimization.
We focus on enhancing our data-driven model’s ability to predict the long-term behavior of ENSO over timescales beyond one year. With the optimal hyper-parameter settings of the realtime filter and the reservoir computing model, the average all-season correlation skill remains above 0.5 for 29 months (Fig. 5). Even discounting the fact that the filtered time series to be modeled is shifted about 5 months into the future due to the asymmetry of the filter function, the performance of the current model succeeds in predicting the dynamics of ENSO for 24 months (i.e., two years). The two-year prediction horizon achieved in this study surpasses the 21 months reported in Hassanibesheli et al. (2022)Hassanibesheli, Kurths, and Boers (2022) using reservoir computing, despite we strictly avoid the use of future information. Furthermore, our result is comparable to the two-year prediction horizon obtained with a deep neural network model Wang et al. (2024), while it should be noted that the target of prediction in this study is the filtered time series.
Figure 5: Prediction skill of the data driven model. All-season correlation skill for ecch forecast lead time is obtained by averaging over 180 predictions between January 2001 and December 2015 based on Eq. (7). The horizontal red dashed line denotes a correlation of 0.5, displayed as a benchmark for the model’s predictive capability.
Figure 6: Comparison between the predicted and the target time series of filtered realtime SST anomaly. (a) 180 predictions and (b) corresponding target. Each row in the two-dimensional plot corresponds to an individual forecast. The horizontal axis indicates the lead time from the start of the prediction, while the vertical axis shows the prediction start time step, with January 2001 set as zero. The maximum forecast lead time is 36 months.
The predicted time series are displayed along with the target in Fig. 6. In several successful cases (e.g. start month 85 and 110), we have been able to capture the major positive and negative peaks of the realtime SST anomaly up to three years in advance.
IV Discussions and conclusions
IV.1 Discussions
In complex phenomena, such as atmosphere-ocean coupled dynamics, multiple processes operating at various speeds are intertwined, forming cross-scale interactions. In such cases it is known that the prediction accuracy of the reservoir computing model varies depending on the passband of the filterHassanibesheli, Kurths, and Boers (2022); Huang, Yang, and Fu (2020). Figure 7 shows the power spectrum of the time series before and after applying the filter. It can be observed that the filter significantly reduce fluctuations with periods shorter than approximately three years and those longer than approximately eight years. For a practical methodology for prediction, it is necessary to identify filter parameter settings that maximize the model’s predictive performance while minimizing the deviation of the filtered time series from its original.
Figure 7: Temporal characteristics of the original and filtered realtime SST anomaly. (a) Power spectrum and (b) autocorrelation. x axis of (a) indicates the frequency in cycles per month
In this paper, fluctuations with periods shorter than one to two years are removed. It is suggested that the reservoir computing model can more easily capture the primary timescale variations of ENSO through this filtering process. By applying the filter, the local minimum at about 22 months and the local maximum at about 44 months of the autocorrelation of the realtime SST anomaly are amplified (Figure 7b). This implies that the periodic fluctuations of about 44 months are relatively strengthened, which in turn enhances the model’s predictive performance on the lead time of half of the period, namely, approximately 22 months.
One characteristic observed in the all-season correlation skill of the reservoir model’s predictions is that, it decreases from unity until reaching a local minimum at around 15 months, and then increases again up to a lead time of 23 months (Figure 5). Similar temporary increase in correlation or decrease in relative error have been reported in previous studies using band-pass filtersHassanibesheli, Kurths, and Boers (2022); Huang and Fu (2023). More detailed investigation into the relationship between the prediction skills and the passband settings of the filter is required in our future study.
IV.2 Conclusions
In this study, using ENSO as an example, we proposed a machine learning methodology that enables long-term prediction of time series in complex phenomena characterized by high-dimensional dynamical systems. By applying a realtime filter that passes temporal scales characteristic of the target time series, and using delayed coordinates of the filtered time series as input variables to a reservoir computing model, we demonstrated improvements in performance of our prediction. A notable feature of this method is that it can be carried out from time series data generation to the pre-forecast process without incorporating any future information, thereby enabling operational realtime prediction. Since the workflow incorporates a Bayesian estimation algorithm to discover the optimal combination of hyperparameters for both the filter and the data-driven model, it suggests that this approach can be applied to dynamical phenomena other than ENSO as well.
Acknowledgements.
TJ was supported by University of Toyama Data Science Promotion Project. Research of TY was partly supported by the JSPS Grants-in-Aid for Scientific Research 24H00186. TM was supported by the JST Moonshot R&D (JPMJMS2389). KN was supported by JSPS KAKENHI Grant No.22K17965. YS was supported by JSPS KAKENHI Grant No.19KK0067, 21K18584, 23H04465 and 24K00537.
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Takuya Jinno: Conceptualization (lead); Data Curation (lead); Formal analysis (lead); Software (equal); Validation (equal); Visualization (lead); Writing - original draft (equal). Takahito Mitsui: Data Curation (supporting); Methodology (supporting); Writing - original draft (supporting). Kengo Nakai: Methodology (lead); Software (supporting); Writing - original draft (supporting). Yoshitaka Saiki: Funding (supporting); Methodology (supporting); Writing - original draft (supporting). Tsuyoshi Yoneda: Conceptualization (supporting); Formal analysis (supporting); Funding acquisition (lead); Software (equal); Validation (equal); Writing - original draft (equal).
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.